OpenStudy (anonymous):

which property is show by 2 + 5 = 5 + 2?

terenzreignz (terenzreignz):

I think I could let this conceptual question off... The fact that 'addition' doesn't care what order the addends are in... 4+9 = 9+4 = 13 is called the 'Commutative Property' (of Addition)